a 20% dye solution is to be mixed with 60% dye solution to get 240L of a 50% solution. how many liters of the 20% and 60% will be needed?
Found 2 solutions by stanbon, ikleyn:
Answer by stanbon(75887)

a 20% dye solution is to be mixed with 60% dye solution to get 240L of a 50% solution. how many liters of the 20% and 60% will be needed?
-------
Equation:
active + active = active
------
0.20x + 0.60(240-x) = 0.50*240
------
20x + 60*240 - 60x = 50*240
------
-40x = -10*240
x = 240/4 = 60 L (amt. of 20% solution needed)
240-x = 180 L (amt. of 60% solution needed)
------------
Cheers,
Stan H.
----------

Answer by ikleyn(52776)

.
For solution of similar problems see the lessons
    - Solving typical word problems on mixtures for solutions   and
    - Mixture problems
in this site.